# TalkPlayData 2: An Agentic Synthetic Data Pipeline for Multimodal Conversational Music Recommendation

### v1

@@ Abstract
We present TalkPlayData 2, a synthetic dataset for multimodal conversational music recommendation generated by an agentic data pipeline. In TalkPlayData 2 pipeline, multiple large language model (LLM) agents are created under various roles with specialized prompts and access to different parts of information, and the chat data is acquired by logging the conversation between the Listener LLM and the Recsys LLM. To cover various conversation scenarios, for each conversation, the Listener LLM is conditioned on a finetuned conversation goal. Finally, all the LLMs are multimodal with audio and images, allowing a simulation of multimodal recommendation and conversation. In the LLM-as-a-judge and subjective evaluation experiments, TalkPlayData 2 achieved the proposed goal in various aspects related to training a generative recommendation model for music. TalkPlayData 2 and its generation code are open-sourced at https://talkpl.ai/talkplaydata2.html.
@@ 2.1 Grounding Music Data, $x_{music}$
To achieve factual data generation, TalkPlayData 2 is primarily based on a source dataset for the details about music items, $x_{music}$. In other words, $y_{music}$, the recommended tracks, must be merely a result of sequential selections of $x_{music}$, the recommendation pool. More details about the source data are provided in subsection 3.1. The music data $x_{music}$ is a set of loosely relevant music items, e.g., music tracks in a listening session (or a playlist as done in (Doh et al. (2025))) in this paper. It is provided with a rich set of metadata, tags, lyrics, audio, and images.
@@ 2.2 Data Generation Agents, $f=\{g_{goal},g_{profile},g_{listener},g_{recsys}\}$
The creation system $f$ is formed by multiple LLM sessions, i.e. a set of agents, each of which has a role – as a listener profiler (Profile LLM, $g_{profile}$), a conversation goal setter (Goal LLM, $g_{goal}$), a listener (Listener LLM, $g_{listener}$), and a recommendation system (Recsys LLM, $g_{recsys}$). This agentic approach presents several advantages compared to relying on a single LLM session as done in (Doh et al. (2025)). Most critically, it systematically prevents the agent from “cheating” by looking at the data provided to other roles. This makes the data generation process highly realistic. For example, a conversation goal is shared with the Listener LLM so that it generates relevant queries to the Recsys LLM and achieves its goal. However, the goal is not shared with the Recsys LLM, whose role is to guess the goal of the Listener LLM through conversation. This approach also provides the role, task, and behavior instructions to each LLM with high clarity, leading to generate high-quality data.
@@ 2.3 The Utilized Capabilities of LLMs, $\theta$
Domain Knowledge Key, chord, and tempo (Zhou et al. (2024); Li et al. (2024))
@@ 2.4 User Profile and Conversation Goal, $x_{profile}$ and $x_{goal}$
Although the LLM generation process is often stochastic, it is well-known that naively sampling multiple times does not lead to diversifying the generation outcomes. Rather, a mode collapse often occurs, where the generated texts become too similar in style and logic Wang et al. (2025); Chen et al. (2025). We observed this in our preliminary experiment - even when using different music items, the generated conversations had very similar styles. To address this issue, we utilize user demographic information and pre-defined conversation goals to generate diverse conversations. Furthermore, utilizing Goal LLM ($g_{goal}$) and Profile LLM ($g_{profile}$), we enhance $x_{profile}$ and $x_{goal}$ to be more suitable for recommending music $x_{music}$, enabling the generation of more diverse and realistic conversations. The details are provided in subsection 3.2.
@@ 3.1 Overview
The foundation of TalkPlayData 2 is the LFM-2b dataset (Schedl et al. (2022)), which provides session-based music listening history data of over 120,000 users spanning more than 15 years (February 2005 to March 2020). Beyond basic metadata (track name, album name, artist name), LFM-2b provides rich additional information including user demographic data (country, gender, age), last.fm genre/style annotations, and ID mappings to Spotify track identifiers. Additional multimodal information is acquired through the provided Spotify track identifiers and the API – preview audio snippets, album art images, release dates, and popularity metrics (as of 2025 July). Finally, we used pretrained music information retrieval models to estimate rich information: Madmom (Böck et al. (2016)) for tempo, key, and chords as well as Whisper (Radford et al. (2023)) for lyrics.
@@
1:Listening sessions dataset $S$, Set of tracks $T$, User Profile $P$, Conversation Goal $G$
2:for each listening session $s$ with $|s|\geq 21$ tracks do
3:  $T_{profile}\leftarrow$ sample 5 tracks of $s$
4:  $T_{pool}\leftarrow$ sample 16-32 tracks of $s$- $T_{profile}$
5:  $P_{base}\leftarrow$ user demographic information of $s$ (age, country, gender)
@@
7:  $P_{final}\leftarrow$ ListenerProfileLLM($T_{profile}$, $P_{base}$)
8:  $G_{final}\leftarrow$ ConversationGoalLLM($G_{base}$, $T_{pool}$)
9:  $Q_{1}\leftarrow$ ListenerLLM($P_{final}$, $G_{final}$, $T_{profile}$) $\triangleright$ First query message ($Q_{1}$)
10:  $M_{1},R_{1}\leftarrow$ RecsysLLM($P_{final}$, $T_{pool}$, $Q_{1}$) $\triangleright$ First music ($M_{1}$), response ($R_{1}$)
@@
13:   $Q_{t}\leftarrow$ ListenerLLM($P_{final}$, $G_{final}$, $T_{profile}$, conversationconversation) $\triangleright$ Listener turn
@@
15:   $M_{t},R_{t}\leftarrow$ RecsysLLM($P_{final}$, $T_{pool}$, conversationconversation) $\triangleright$ Recsys turn
@@ 3.2 Generation Process
The overall process iterates over the listening sessions $S$ and converts each session $s\in S$ into a conversation. Each session $s$ consists of a list of tracks $T$ and basic user demographic information $P$, including age group, gender, and country. For each conversation, we require sessions containing at least 21 tracks to ensure a sufficient recommendation pool. From each session, 5 tracks are sampled as profiling tracks ($T_{profile}$) to inform the conversation style sampling, while another 16-32 tracks form the recommendation pool ($T_{pool}$).
For a single conversation, four separate LLM sessions are created, each of which corresponds to $_{profile}, gg$$_{goal}$, $g_{listener}$, and $g_{recsys}$, respectively. Their instructions are carefully designed after many iterations of reviews and updates to ensure correct behaviors and response formats. As outlined in Algorithm 1 and Figure 1, the overall generation process consists of two stages: i) profiling and goal generation, and ii) conversation generation. During the first stage (lines 1-7), we create a user profile $P_{final}$ based on profiling tracks $T_{profile}$ and demographic information $P_{base}$, and a conversation goal $G_{final}$ from base goal templates $G_{base}$. This adaptive customization is responsible for steering the generated conversation in various styles. The conversation generation stage (lines 8-14) follows with alternating API calls between the listener and the recommendation systems, where queries $Q_{t}$, music recommendations $M_{t}$, and responses $R_{t}$ build upon the conversation history while maintaining coherence with the conversation goals $G_{final}$. The detailed instructions and responses for the LLMs are provided in Appendix B.
@@
The role of the Listener Profile LLM is to analyze the profiling tracks and infer high-level preference information of the listener, given factual demographic information, such as gender, age group, and country. The LLM combines the provided demographic profile with the track analysis to estimate musical preferences including preferred musical culture, top artist, and top genre. All the text data (metadata, tags, and lyrics), audio, and image data described in subsection 3.1 are provided to the LLM.
@@
The Conversation Goal defines the session-level goal that the listener wants to achieve through conversation with the recommendation system. To guarantee the overall diversity of the conversations in TalkPlayData 2, a diverse set of conversation goals is needed; while each conversation goal should be plausible given the recommendation pool. Before the generation process, a set of 44 conversation goal templates is prepared. A template is defined by two properties, the topic and the specificities. In total, 11 topics are defined to cover various types of multi-modal music discovery conversations, as listed in Table 3. These topics decide which aspect of music the conversation will be based on. The specificities define how specific the query and the target music are, resulting in 4 cases as in Table 3
@@ I Cultural & Geographic “Music from Alaska”
Based on the profiling tracks, the conversation goal, and the listener profile, the conversation is initiated by the Listener LLM ($g_{listener}$). On Recsys LLM ($g_{recsys}$), after being initialized with the listener profile and the recommendation pool (but not the conversation goal), it starts to respond to the Listener LLM’s initial query. In the subsequent turns (from Turn 2), the Listener LLM actively engages with the recommended music by listening to the audio samples and seeing the album artwork before formulating responses. This multimodal interaction allows the Listener LLM to provide more nuanced and informed feedback about the recommendations, to which Recsys LLM then makes subsequent recommendations. Finally, in every turn, both the Listener and the Recsys LLMs generate ‘thought’ before generating their response message to each other. A ‘thought’ is used to analyze the input message (from the other LLM).
@@ 4.1 Statistics
Table 4 summarizes the basic statistics of TalkPlayData 2. It consists of 1,000 conversations in the test split, where the cold and the warm user sets are defined by if the user id exists in the train split.  [^1] The data generation is work in progress and on track to have generated over 20000 conversations by end of 2025 September. This manuscript will be updated accordingly. The thought is significantly longer than the messages, demonstrating the dataset’s emphasis on chain-of-thought reasoning.
Table 4: Dataset Statistics of TalkPlayData 2
Train Test
Metric Train All Test All Test Warm Test Cold
Total Conversation TBD 1000 800 200
Unique Users TBD 500 400 100
Unique Tracks TBD 6744 5562 1483
Unique Artists TBD 2998 2521 838
Vocabulary Size (words) TBD 16106 14570 7586
Average Message Length (words) TBD 42.3 42.0 43.4
Average Thought Length (words) TBD 102.3 101.2 107.1
Figure 2: User Distribution of TalkPlay2 Test Dataset
Figure 2 shows the user demographic distribution of the TalkPlayData 2 test set split. The test set includes 500 users from 53 different countries. When visualizing the distribution of users across continents, we observe that Europe and the Americas are dominant despite balanced sampling efforts. In terms of age distribution, teenagers and the 20s constitute the majority of users. Gender distribution shows a nearly uniform pattern across categories.
@@ 4.3 Human Evaluation
Table 6: Comparison of conversational music recommendation datasets. Type stands for subject of conversation. Relevance and Naturalness show Mean Opinion Score of 5 Likert Scale.
@@
As shown in Table 6, TalkPlayData 2 achieves the highest scores in both dimensions. The high relevance score demonstrates the effectiveness of our multimodal approach, where LLMs consider both audio and visual aspects of music during recommendation, leading to more accurate and contextually appropriate suggestions compared to text-only approaches (Doh et al. (2024; 2025)). The strong naturalness score highlights the effectiveness of our multi-LLM framework. By orchestrating interaction between the Conversation Goal LLM and the Profile LLM, the system enables naturalistic exchanges between the Listener LLM and the RecSys LLM, thereby improving both conversational coherence and user simulation. These results suggest that our approach of using multiple LLMs creates more engaging and effective conversational recommendations compared to both human conversations Chaganty et al. (2023) and single-LLM approaches (Doh et al. (2024; 2025)).

### v4

@@ Abstract
We present TalkPlayData 2, a synthetic dataset for multimodal conversational music recommendation generated by an agentic data pipeline. In the proposed pipeline, multiple large language model (LLM) agents are created under various roles with specialized prompts and access to different parts of information, and the chat data is acquired by logging the conversation between the Listener LLM and the Recsys LLM. To cover various conversation scenarios, for each conversation, the Listener LLM is conditioned on a finetuned conversation goal. Finally, all the LLMs are multimodal with audio and images, allowing a simulation of multimodal recommendation and conversation. In the LLM-as-a-judge and subjective evaluation experiments, TalkPlayData 2 achieved the proposed goal in various aspects related to training a generative recommendation model for music. TalkPlayData 2 and its generation code are open-sourced at https://talkpl.ai/talkplaydata2.
@@ 2.1 Grounding Music Data, $x_{music}$
To achieve factual data generation, TalkPlayData 2 is primarily based on a source dataset for the details about music items, $x_{music}$. In other words, $y_{music}$, the recommended tracks, must be merely a result of sequential selections of $x_{music}$, the recommendation pool. More details about the source data are provided in subsection 3.1. The music data $x_{music}$ is a set of loosely relevant music items, e.g., music tracks in a listening session in this paper (equivalent to playlists as in Doh et al. (2025)). It is provided with a rich set of metadata, tags, lyrics, audio, and images.
@@ 2.2 Data Generation Agents, $f=\{g_{goal},g_{profile},g_{listener},g_{recsys}\}$
The creation system $f$ is formed by multiple LLM sessions, i.e., a set of agents, each of which has a role – as a listener profiler (Profile LLM, $g_{profile}$), a conversation goal setter (Goal LLM, $g_{goal}$), a listener (Listener LLM, $g_{listener}$), and a recommendation system (Recsys LLM, $g_{recsys}$). This agentic approach presents several advantages compared to relying on a single LLM session such as Doh et al. (2025). Most critically, it systematically prevents the agent from ‘cheating’ by looking at the data provided to other roles. This makes the data generation process highly realistic. For example, a conversation goal is shared with the Listener LLM so that it generates relevant queries to the Recsys LLM and achieves its goal. However, the goal is not shared with the Recsys LLM, whose role is to guess the goal of the Listener LLM through conversation. This approach also provides the role, task, and behavior instructions to each LLM with high clarity, leading to generate high-quality data.
@@ 2.3 The Utilized Capabilities of LLMs, $\theta$
Domain Knowledge Key, chord, and tempo (Zhou et al. (2024); Li et al. (2024b))
@@ 2.4 User Profile and Conversation Goal, $x_{profile}$ and $x_{goal}$
Although the LLM generation process is often stochastic, it is well-known that naively sampling multiple times does not lead to diversifying the generation outcomes. Rather, a mode collapse often occurs, where the generated texts become too similar to each other in style and logic (Wang et al. (2025); Chen et al. (2025)). We observed this in our preliminary experiment - even when using different music items, the generated conversations had very similar styles. To address this issue, we utilize user demographic information and pre-defined conversation goals to generate diverse conversations. Furthermore, utilizing Goal LLM and Profile LLM, we enhance $x_{profile}$ and $x_{goal}$ to be more suitable for recommending music $x_{music}$, enabling the generation of more diverse and realistic conversations. The details are provided in subsection 3.2.
@@ 3.1 Overview
The foundation of TalkPlayData 2 is the LFM-2b dataset (Schedl et al. (2022)), which provides session-based music listening history data of over 120,000 users spanning more than 15 years (February 2005 to March 2020). Beyond basic metadata (track, album, artist name), LFM-2b provides rich additional information including user demographic data (country, gender, age), last.fm genre/style annotations, and ID mappings to Spotify track identifiers. Additional multimodal information is acquired through the provided Spotify track identifiers and the API – preview audio snippets, album art images, release dates, and popularity metrics (as of 2025 July). Finally, we used pretrained music information retrieval models to estimate rich information: Madmom (Böck et al. (2016)) for tempo, key, and chords as well as Whisper (Radford et al. (2023)) for lyrics.
@@
1:Listening sessions $S$, Set of tracks $M$, User Profile $U$, Conversation Goal $G$, Query message $Q$, Recommended Music $M_{t}$, Response $R$, Thought $T$, Progress Towards Goal $P$
2:for each listening session $s$ with $|M|\geq 21$ tracks do
3:  $M_{profile}\leftarrow$ sample 5 tracks of $M$
4:  $M_{pool}\leftarrow$ sample 16-32 tracks of $M$- $M_{profile}$
5:  $U_{base}\leftarrow$ user demographic information of $s$ (age, country, gender)
@@
7:  $U_{final}\leftarrow$ ListenerProfileLLM($U_{profile}$, $U_{base}$)
8:  $G_{final}\leftarrow$ ConversationGoalLLM($G_{base}$, $M_{pool}$, $U_{final}$)
9:  $Q_{1}\leftarrow$ ListenerLLM($U_{final}$, $G_{final}$, $M_{profile}$) $\triangleright$ First query message ($Q_{1}$)
10:  $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{T^{t}_{1}}$, $M_{1},R_{1}\leftarrow$ RecsysLLM($P_{final}$, $T_{pool}$, $Q_{1}$) $\triangleright$ First track ($M_{1}$), response ($R_{1}$)
@@
13:   $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{P_{t},T^{l}_{t}}$, $Q_{t}\leftarrow$ ListenerLLM($U_{final}$, $G_{final}$, $M_{profile}$, conversationconversation)
@@
15:   $\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}{T^{r}_{t}}$, $M_{t},R_{t}\leftarrow$ RecsysLLM($U_{final}$, $M_{pool}$, conversationconversation)
@@ 3.2 Generation Process
The overall process iterates over the listening sessions $S$ and converts each session $s\in S$ into a conversation. Each session $s$ consists of a list of tracks $M$ and basic user demographic information $U$, including age group, gender, and country. For each conversation, we require sessions containing at least 21 tracks to ensure a sufficient recommendation pool. From each session, 5 tracks are sampled as profiling tracks ($M_{profile}$) to inform the conversation style sampling, while another 16-32 tracks form the recommendation pool ($M_{pool}$).
For a single conversation, four separate LLM sessions are created, each of which corresponds to $_{profile}, gg$$_{goal}$, $g_{listener}$, and $g_{recsys}$, respectively. Their instructions are carefully designed after many iterations of reviews and updates to ensure correct behaviors and response formats. As outlined in Algorithm 1 and Figure 1, the overall generation process consists of two stages: i) profiling and goal generation, and ii) conversation generation. During the first stage (lines 1-8), we create a user profile $U_{final}$ based on profiling tracks $M_{profile}$ and demographic information $U_{base}$; and a conversation goal $G_{final}$ from base goal templates $G_{base}$, recommendation pool $M_{pool}$, and user profile $U_{final}$. This customization is responsible for diversifying the conversation in style. The conversation generation stage (lines 11-15) follows with alternating API calls between the listener and the recommendation systems, where queries $Q_{t}$, music recommendations $M_{t}$, and responses $R_{t}$ build upon the conversation history while maintaining coherence with the conversation goals $G_{final}$. The detailed instructions and responses for the LLMs are provided in Appendix B.
@@
The role of the Listener Profile LLM is to analyze the profiling tracks and infer high-level preference information of the listener, given factual demographic information, such as gender, age group, and country. During generating TalkPlayData 2 as well as using it, this profile information provides the personalization aspect, which is crucial in modern recommendation systems and music consumption (Schedl et al. (2021); Kaminskas & Ricci (2012); North & Hargreaves (2008)). The LLM combines the provided demographic profile with the track analysis to estimate musical preferences including preferred musical culture, top artist, and top genre. All the text data (metadata, tags, and lyrics), audio, and image data described in subsection 3.1 are provided to the LLM.
@@
The Conversation Goal defines the session-level goal that the listener wants to achieve through conversation with the recommendation system. To guarantee the overall diversity of the conversations in TalkPlayData 2, a diverse set of conversation goals is needed; while each conversation goal should be plausible given the recommendation pool (Li et al. (2024a)). Before the generation process, a set of 44 conversation goal templates is prepared. A template is defined by two properties, the topic and the specificities. In total, 11 topics are defined to cover various types of multi-modal music discovery conversations, as listed in Table 3. These topics decide which aspect of music the conversation will be based on and cover recommendation scenarios based on audio (Deldjoo et al. (2024); Van den Oord et al. (2013), lyrics (Patra et al. (2017); Vystrčilová & Peška (2020)), visual information (Saito & Itoh (2011); Libeks & Turnbull (2011)), and emotion (Han et al. (2010)). The specificities define how specific the query and the target music are, resulting in 4 cases as in Table 3. This two-dimensional formalization of LL, HL, LH, and HH provides a structured view of recommendation scenarios such as exploratory search (Marchionini (2006); Schedl et al. (2015)), lookup tasks (Marchionini (2006)), and query granularity (Sun & Zhang (2018); Jannach et al. (2021).
@@ I Cultural & Geographic “Music from Alaska”
Based on the profiling tracks, the conversation goal, and the listener profile, the conversation is initiated by the Listener LLM ($g_{listener}$). On Recsys LLM ($g_{recsys}$), after being initialized with the listener profile and the recommendation pool (but not the conversation goal), it starts to respond to the Listener LLM’s initial query. In the subsequent turns (from Turn 2), the Listener LLM actively engages with the recommended music by listening to the audio samples and seeing the album artwork before formulating responses. This multimodal interaction allows the Listener LLM to provide more nuanced and informed feedback about the recommendations, to which Recsys LLM then makes subsequent recommendations. The Listener LLM also labels whether the Recsys LLM’s recommendation is making positive progress towards achieving the goal. This information is expected to be used as a ‘preference’ signal during training recommendation models. Finally, in every turn, both the Listener and the Recsys LLMs generate ‘thought’ before generating their response message to each other. A ‘thought’ is used to analyze the input message (from the other LLM), increasing interpretability during both dataset creation and utilization.
@@ 4.1 Statistics
Table 4: Comparison among Conversational Music Recommendation Datasets. Gray text indicates Closed-Source Datasets. CS refers to Cold-Start Split.
Dataset Profile Goal Thought CS Conv. Track User Turns
JAMSessions N/A 112K 100k 104K 1.00
Text2Track N/A 1M 500K - 1.00
CPCD 0.1K 107K - 5.70
LP-MusicDialog 288K 391K - 4.97
TalkPlayData 1 532K 406K - 6.95
TalkPlayData 2 (Ours) 16.2K 47K 9k 8.00
Table 5: Statistics of the TalkPlayData 2
Statistics Training Evaluation
# of Conversations 15199 1000
# of Warm Users - 371
# of Cold Users - 129
# of Total Users 8591 500
# of Warm Tracks - 3779
# of Cold Tracks - 2982
# of Total Tracks 43597 6761
Avg. Query Length 44.7 39.8
Avg. Response Length 36.6 44.6
Avg. Thought Length 89.3 101.7
Table 4 presents a comparison among conversational music recommendation datasets. We compare against two closed-source single-turn conversation datasets (Palumbo et al., 2024; Melchiorre et al., 2025), a human conversation dataset (Chaganty et al., 2023), and two LLM-based synthetic datasets (Doh et al., 2024; 2025). While TalkPlayData 2 does not have a large number of conversations, it represents the dataset most similar to real music recommendation scenarios, featuring 1) long conversation turns, 2) user profiles, 3) conversation goals, 4) chain-of-thought, and 5) cold-start splits.
Table 5 summarizes the key statistics of TalkPlayData 2, which consists of a 15.5K training set and 1K test set divided using chronological splitting to reflect real-world deployment scenarios. The test set includes 100 cold users and 6,507 cold tracks for evaluating cold-start scenarios, while a distinctive characteristic is the comprehensive inclusion of user queries, assistant responses, and detailed thought processes (averaging 89.4-102.3 words) that provide valuable insights into the reasoning behind preferences and recommendations, enabling interpretable music recommendation systems.
@@ 4.3 Human Evaluation
Table 7: Comparison of conversational music recommendation datasets. Type stands for the subject of conversation. Relevance and Naturalness show Mean Opinion Scores of 5 Likert Scale.
@@
As shown in Table 7, TalkPlayData 2 achieves the highest scores in both dimensions. The high relevance score demonstrates the effectiveness of our multimodal approach, where LLMs consider both audio and visual aspects of music during recommendation, leading to more accurate and contextually appropriate suggestions compared to text-only approaches (Doh et al. (2024; 2025)). The strong naturalness score highlights the effectiveness of our multi-LLM framework. By orchestrating interaction between the Conversation Goal LLM and the Profile LLM, the system enables naturalistic exchanges between the Listener LLM and the RecSys LLM, thereby improving both conversational coherence and user simulation. These results suggest that our approach of using multiple LLMs creates more engaging and effective conversational recommendations compared to both human conversations (Chaganty et al. (2023)) and single-LLM approaches (Doh et al. (2024; 2025)).
4.4 Ablation Study
The goal of the ablation study is to provide empirical evidence that analyzes and supports the design choices of TalkPlayData 2. The experiments are conducted with three configurations: removing the conversation goal (A1), the listener Profile (A2), and both (A3). In each configuration, 50 conversations are generated in total, all based on the same base data subset of LFM-2b. Then, an LLM judge is prompted to classify each conversation into one of the 4 specificities and 11 topics. We use i) the Kullback–Leibler divergence (KLD) to measure how close the empirical distributions are to a uniform distribution and ii) Coverage, the proportion of classes that have non-zero items, to track any potential strong category biases.
As presented in Table 8, only the proposed pipeline achieved the full coverage on Specificity and Topic, as well as the lowest $KLD_{u}$ in Specificity and the second lowest $KLD_{u}$ in Topic. In detail, first, the overall importance of the goal is clear, based on the significant degradation from every aspect in A1 and A3. This is expected, since the goal is the only prompt where the Specificity and the Topic are defined. Second, the impact of the profile is more nuanced, since in A2, the metrics on the Topic do not indicate any issues, while the diversity of the Specificity shows severe regression, i.e. the goal alone is enough to generate conversations with diverse topics, but not with diverse specificities. We conjecture this is due to the close relationship between the Specificity and listener behaviors, as well as the sequential order that the profile conditions the goal. The profile includes not only demographics but also open-vocabulary attributes such as preferred musical culture, top artists, and genres – altogether, seemingly contributing to the diversity of how specific a user would query and expect; and when there is a lack of such information, the Goal LLM is biased towards certain specificities.
Table 8: KL divergence to uniform (KLDu, $\downarrow$) and coverage ($\uparrow$) across ablations.
KLDu (Specificity) KLDu (Topic) Coverage (Specificity) Coverage (Topic)
TalkPlayData 2 0.240 0.110 1.000 1.000
A1: no goal 0.316 0.700 0.750 0.455
A2: no profile 0.553 0.045 0.750 1.000
A3: no goal+profile 0.395 0.451 0.750 0.455
5 Related Works
Conversational Music Recommendation: Conversational recommendation systems provide recommendations through natural language dialog with users, requiring both multi-turn recommendation capabilities and natural language response generation (Goker & Thompson, 2000; Christakopoulou et al., 2016; Zhang et al., 2018). These systems predict ranked lists of relevant music tracks based on conversation history and user queries while understanding preferences, context, and diverse query types. Beyond recommendation, systems generate engaging responses that describe recommendations and assist users in music exploration through natural language explanations (Doh et al., 2024).
In the music domain, early approaches leveraged dense embeddings to find appropriate music by computing similarity between multi-turn history embeddings and item embeddings (Chaganty et al., 2023; Doh et al., 2024; Melchiorre et al., 2025), but these methods suffered from limited response generation capabilities. Recent studies have employed LLMs to enable both multi-turn retrieval through semantic token generation and natural language response generation through text token generation (Palumbo et al., 2024; Doh et al., 2025). However, these approaches have been constrained by data quality issue, resulting in conversations with short turns, insufficient user profiles, and difficulties in training on diverse conversational topics (more detail in Table 4).
LLM-based Dataset Generation: Recent studies have shown that LLMs can generate large, useful corpora for conversational music recommendation tasks. LP-MusicDialog (Doh et al., 2024) and Talk the Walk (Leszczynski et al., 2023) synthesize hundreds of thousands to over a million user-item conversations by grounding LLM utterances in curated music collections and attribute filtering. More recently, TalkPlay (Doh et al., 2025) also used a large-scale dataset synthesized by LLM for generative recommendation of music. These works demonstrate that synthetic conversational dataset can be used to train and evaluate music retrieval and dialogue models, but they typically rely on single-session LLM generation or primarily text grounding.
Recent research also emphasizes multi-agent pipelines to improve realism, diversity, and safety. Multi-agent systems such as The Fellowship of the LLMs (Arif et al., 2024), MAG-V (Sengupta et al., 2024), and AgentSGEN (Xuan et al., 2025) show how separating roles (generator, judge, editor) or adding verification loops reduces mode collapse and improves semantic fidelity. SampleLLM (Gao, 2025) targets distribution alignment for tabular recommendation data, while recent surveys (Nadas, 2025) summarize key techniques and pitfalls for LLM-driven synthesis. Together, these works illustrate the growing importance of synthetic data generation for building conversational recommender systems successfully.